We have monitored the affected services and all looks stable now. As such we are closing this incident.
Nov 15, 01:11 GMT
We have monitored the services and it remains stable. We will continue to monitor the device to ensure stability.
Nov 14, 05:02 GMT
We have added the module back and all services are back to normal. We will continue to monitor the device to ensure stability.
Nov 14, 01:07 GMT
The impacted module has been rebooted and behavior seems to be normal. We will add the module back to service tonight. We are further coordinating with vendor to identify the root cause.
Nov 13, 17:09 GMT
Our on-site engineer has removed the faulty module from the switch. We are coordinating with the vendor for a permanent fix.
Nov 13, 14:31 GMT
Our engineer is now at datacenter and is further troubleshooting the issue in cordination with the vendor. Further updates will be provided in due course.
Nov 13, 13:07 GMT
Our engineer will visit the datacenter today to check the device fault and we will further coordinate with the vendor. Further updates will be provided in due course.
Nov 13, 06:23 GMT
Our engineers are actively coordinating with the vendor to get it resolved soon. Further updates will be provided in due course.
Nov 12, 22:53 GMT
We have identified a faulty module in one of our core switches. Resiliency is affected for firewall traffic. Due to this customers may experience a performance degradation. We have our engineers actively working on it.
Nov 12, 16:53 GMT